OpenStudy (anonymous):

Solve the inequality: x^2-3x-18<0

OpenStudy (sirm3d):

factor the left side of the inequality

OpenStudy (anonymous):

oh and?...

OpenStudy (anonymous):

x^2 - 3x - 18 < 0 x^2 - 3x < 18 x^2 - x > - 6 * you would switch the inequality sign because you divided by a negative. x > 12 I'm pretty sure that would be the answer, I'm not 100% though.

OpenStudy (sirm3d):

here's the detail \[x^{2}-3x-18<0\] \[\left( x+3 \right)\left( x-6 \right)<0\]

OpenStudy (anonymous):

\[x^2 - 3x -18 < 0\] \[(x-6)(x+3)\] This would lead to: \[x-6 < 0\] or \[x+3<0\] Which means that either \[x<6\] or\[x<-3\] Treat the inequality as if it were an equation but be mindful that by multiplying or dividing by minus signs changes the direction of the inequality sign

OpenStudy (sirm3d):

the roots of the left hand side are -3 and 6 choose a value in each intervals \[x<-3\] \[-3<x<6\] \[6<x\]

OpenStudy (sirm3d):

comment on Muli97 solution if \[\left( x+3 \right)\left( x-6 \right)<0\] then one factor is positive and the other factor is negative